Weekly JLC Physics Group Meeting

Place: BLDG 3, KEK
Time:  10:30, 97/02/07

[0] Agenda
   0.1) Kakenhi Report
   0.2) Workstation Maintenance
   0.3) CAL Subgroup Status
   0.4) CDC Subgroup Status
[1] Kakenhi Report (Miyamoto)
   Most parts are now available.
   Those who are responsible for the missing parts should
   work harder!

[2] Workstation Maintenance

   Should decide the 1997 budget for workstation maintenance.
   If we follow the 1996 example, it would cost 3.6Myen for
   	jlcux1                                : hardware + software
   	jlcux2, www-jlc, jlcsv1, jlcuxm, kobe : hardware only.
   We would have to include jlcuxf and jlcuxjk, which causes
   additional expenses.
   Decided to try to save as much money as possible, by cutting the
   maintenance contract to minimum possible.
[3] CAL Subgroup Status (Toeda)
   3.1) Test Calorimeter Construction
   	Calorimeter assembly: almost finished
			Is disassembling the calorimeter really
			necessary for transporting to the PS counter
			---> to be discussed again in the presence of
   3.2) Schedule
   	Early Feb.	: light shielding and cosmic ray test
	Feb.21 - 25	: setup in PS
	Feb.27 -	: beam run 
	Mar.17		: finish beam run

[4] CDC Subgroup Status (Yamaguchi)

   4.1) Cosmic Ray Data Status
   	Finish the data taking in the central region for both
   	90- and 45-degree tracks.
   	---> Got lower O2 concentration (50ppm compared to 
   	     usual 100ppm) data with times-2 gas flow.
   	     The pulse height seems bigger.
   	     45-degree data also seem to show a higher pulse
   	     height spectrum.

   4.2) Problem

   	Moved the trigger counter toward the H.V.-side end of
   	the chamber and started taking 90-degree data.
	---> Data taking rate became significantly lower.
	     Under investigation.